OpenStudy (anonymous):

(2^x+ 7)(x -5) -3 (x^2+ 5x)?

OpenStudy (mathstudent55):

First, multiply out the two binomial on the left using FOIL. Then, distribute the -3. Then combine like terms.

OpenStudy (whpalmer4):

Expand it out and collect like terms. Is that really 2^x and not x^2?

OpenStudy (whpalmer4):

Also, when subtracting -3(x^2+5x) don't forget to change the sign on the 5x term as well. That's a frequent mistake! -3(x^2 + 5x) = -3x^2 -15x, NOT -3x^2 + 15x

OpenStudy (whpalmer4):

Yeah, that must be x^2, not 2^x...

OpenStudy (anonymous):

all sides of a triangle are known to be intergers. Two of the sides of the triangle are 8 and 12. How many different values are there for the third side?

OpenStudy (whpalmer4):

Don't put new problems in existing threads, open new questions instead. And give medals if you expect additional help from people who answer your questions :-)

OpenStudy (anonymous):

okay thank you :)

OpenStudy (mathstudent55):

(x^2+ 7)(x -5) -3 (x^2+ 5x) = x^3 -5x^2 + 7x - 35 - 3x^2 - 15x Now collect like terms